how many grams of barium sulfate are produced if 25.34 mL of 0.113M BaCl2 completely react given the reaction: BaCl2(aq)+Na2SO4(aq)=BaSO4(s)+2NaCl(aq)
Answer #1

Consider reaction , BaCl2 (aq) + Na2SO4 (aq)  \rightarrow BaSO4 (s)+2 NaCl (aq)

From reaction, 1 mol  BaCl2 \equiv 1 mol Na2SO4\equiv 1 mol BaSO4  \equiv 2 mol NaCl

i e 1 mole  BaSO4 will be produced from 1 mole  BaCl2.

We have, [  BaCl2 ] = No. of moles of  BaCl2 / volume of solution in L

\therefore No. of moles of  BaCl2 = [  BaCl2 ] \times volume of solution in L

Here we have,  [  BaCl2 ] = 0.113 M and volume of solution = 25.34 ml = 0.02534 ml.

\therefore No. of moles of  BaCl2 = 0.113 mol / L \times 0.02534 L = 0.002863 mol

We have relation, 1 mol  BaCl2 \equiv 1 mol BaSO4

Hence, No. of moles of  BaCl2 reacted = no. of moles of BaSO4 produced = 0.002863 mol

We have , No. of moles = Mass / Molar mass

\therefore Mass = No. of moles \times Molar mass

Molar mass of BaSO4 = 137.34 + 32.06 + ( 4 \times 16.00) = 233.40 g/mol

\therefore Mass BaSO4 produced in the reaction = 0.002863 mol \times 233.40 g /mol = 0.668 g

ANSWER :  Mass BaSO4 produced in the reaction = 0.668 g